Circular No. 5850

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993X IN NGC 2276 R. R. Treffers, A. V. Filippenko, B. Leibundgut, Y. Paik, and L. F. M.

Lee, University of California at Berkeley; and M. W. Richmond, Princeton University, report their discovery of an apparent supernova in NGC 2276 (α = 7^{h}10^{m}31^{s}, δ = +85^{o}50′.9, equinox 1950.0), located ∼ 30′′ east and 69′′ north of the galaxy’s nucleus. The object was found during the Leuschner Observatory Supernova Search, which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. It was visible at mag R = 16.3 ± 0.4 on Aug. 22 UT, with confirmation at about the same brightness on Aug. 23 and 25. It could not be detected to limiting mag R = 17.5-18.0 in images obtained on Mar. 6, Apr. 2, May 1, and 8.

RX J1940.2–1025 S. Rosen, C. Done, and M. Watson, University of Leicester; and G.

Madejski, Goddard Space Flight Center, report: “We have an optical identification for an x-ray source close to the Seyfert galaxy NGC 6814. This source is now believed to be the origin of the periodic signals previously associated with the galaxy (Done et al. 1992, Ap.J. 400, 138). The x-ray source (RX J1940.2–1025), discovered in a recent ROSAT PSPC observation, has an average PSPC count rate of 1.4 counts/s. It lies ∼ 37′ from NGC 6814, but close enough to be in the field-of-view of non-imaging x-ray detectors, and the x-ray data clearly show the 3.4-hr period previously associated with NGC 6814. The ROSAT PSPC spectrum is consistent with it being an AM Her-type star (although other interpretations are possible). Using the ROSAT PSPC position, the x-ray object was identified on Aug. 18 using the Anglo-Australian Telescope (+ RGO spectrograph). The counterpart is an emission-line object of mag 16-17 showing strong Balmer and He II 468.6-nm emission lines. The width of the lines (FWZI up to 3000 km/s), and the large He II 468.6-nm/Hβ line ratio strongly support the AM Her classification of this object. The position of the object determined by astrometry of the POSS plates is α = 19^{h}40^{m}11^{s}.38, δ = −10^{o}25′24′′.5 (equinox 2000.0; estimated error 1′′). The optical light curve of this object reconstructed from the time-resolved spectroscopy shows a strong and complex modulation.”

1993 August 27 Daniel W. E. Green